# Groups which are metrically weakly sofic with respect to word norms

### v1

@@ 3.3 When metric RF does not depend on metric
Assume that $G$ is normally generated by finite $S_{1}$ and $S_{2}$. Let $\parallel\cdot\parallel_{1}$ and $\parallel\cdot\parallel_{2}$ be the corresponding word metrics of $G$. The following question looks very interesting. Does the property that $G$ is metrically LEF (metriccally RF or metrically weakly sofic) with respect to one of the norms $\parallel\cdot\parallel_{1}$ and $\parallel\cdot\parallel_{2}$ imply the same property for another norm? We do not even know if these word norms generate the same topology in the following sense. Note that the topology defined by $\parallel\cdot\parallel_{i}$, $i\in\{1,2\}$, according to the recipe of Section 2.1, is the discrete $\{0,1\}$-topology. On the other hand if we exclude from the sub-base of all $B_{<r}(g)$ the balls of size 1 then we obtain a coarser topology which is not necessarily discrete. We will say that this is the corrected topology defined by the norm. Using Theorem 3.3 we see the following statement.
@@ 4 Metric weak soficity
Assume that the set $S\subseteq\mathsf{F}$ defining the word metric $\parallel\cdot\parallel_{N}$ is finite. The metric group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically weakly sofic if and only if for any finite sequence $g_{1}$, $g_{2}$,…, $g_{k}\in B^{\mathsf{F},\parallel\cdot\parallel}_{1}(1)$ and $h_{1}$, $h_{2}$,…, $h_{l}$ from $N$ (possibly with repetitions) the closure of the product $g_{i_{1}}^{\mathsf{F}}\cdot g_{i_{2}}^{\mathsf{F}}\ldots\cdot g_{i_{m}}^{\mathsf{F}}\cdot h_{1}^{\mathsf{F}}\cdot h_{2}^{\mathsf{F}}\ldots h_{l}^{\mathsf{F}}$ is a subset of $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$.
@@
Necessity. Assume $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$. Let $s=\parallel wN\parallel_{N}$. Thus $s>m$. Let $n$ be a natural number such that
@@
Let $\tilde{\phi}$ be the homomorphism $\mathsf{F}\to H$ defined by $\phi$ on the free basis of $\mathsf{F}$. It is easy to see that for any $v\in B^{\mathsf{F},|\cdot]}_{<n}$ we have $\ell((\tilde{\phi}(v))^{-1}\phi(vN))<2n\varepsilon$ (we should take double $n\varepsilon$ since inverses of the elements of the base can also appear in $v$). It suffices to show that $\tilde{\phi}(w)$ does not belong to
@@
We view $\phi$ on cosets from $B^{\mathsf{F},|\cdot]}(1)N$ as a map which assigns $\phi(f)$ to $fN$ with $f\in B^{\mathsf{F},|\cdot]}(1)$. In order to verify the weak soficity condition for $\phi$, $\varepsilon$ and elements of $FN$ note that when $f_{1}N\cdot f_{2}N=f_{3}N$ for $f_{1},f_{2},f_{3}\in F$ we have that $\phi(f_{1})\phi(f_{2})(\phi(f_{3}))^{-1}$ is equal to some $\phi(h)$ with $h\in N$, i.e an element of the norm $\varepsilon$.
@@
In the assumptions of the theorem admit that $S$ can be infinite and can meet infinitely many conjugacy classes in $\mathsf{F}$. Then if the metric group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically weakly sofic then for any finite sequence $g_{1}$, $g_{2}$,…, $g_{k}\in B^{\mathsf{F},\parallel\cdot\parallel}_{1}(1)$ and $h_{1}$, $h_{2}$,…, $h_{l}$ from $N$ (possibly with repetitions) the closure of the product $g_{i_{1}}^{\mathsf{F}}\cdot g_{i_{2}}^{\mathsf{F}}\ldots\cdot g_{i_{m}}^{\mathsf{F}}\cdot h_{1}^{\mathsf{F}}\cdot h_{2}^{\mathsf{F}}\ldots h_{l}^{\mathsf{F}}$ is a subset of $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$.
@@ 5.1 Characterization of metric LEF
Our characterization of metric LEF is slightly long, but it clarifies the situations of Sections 3 and 4. We preserve the notation of Section 3.1.
@@
Assume that the set $S\subseteq\mathsf{F}$ defining the word metric $\parallel\cdot\parallel_{N}$ is finite. The metric group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically LEF if and only if for any finite set $F\subset\mathsf{F}$ representing pairwise distinct elements modulo $N$, for any natural number $m$ and any $w\in\mathsf{F}\setminus B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ there is a homomorphism $\varphi$ from $\mathsf{F}$ onto a finite group $H$ such that $\varphi(w)\not\in\varphi(B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N)$ and the corresponding map $fN\to\varphi(f)$, $f\in F$, is a partial isomorphism $\mathsf{F}/N\to H$ defined on $\{fN|f\in F\}$.
@@
Necessity. Let $g_{1}$, $g_{2}$,…, $g_{m}$ be a list of the representatives of conjugacy classes of $S^{\mathsf{F}}$. Thus
$B^{\mathsf{F},\parallel\cdot\parallel}_{1}(1)=g_{1}^{\mathsf{F}}\cup g_{2}^{\mathsf{F}}\ldots\cup g_{m}^{\mathsf{F}}.$
Take any finite $F$ as in the formulation of the theorem, and $m$ and $w\in\mathsf{F}\setminus B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$. We may assume that $F$ contains $w$, $g_{1}$, $g_{2}$,…, $g_{m}$ and the free generators of $\mathsf{F}$ and, furthermore, it is closed under the operation of taking subwords of a word. Choose a partial map $\psi$ from $\mathsf{F}/N$ to some finite $(H,\ell)$ which is defined on $\{fN|f\in F\}$ as an isomorphism and
$\ell(\psi(g_{i}N))=\parallel g_{i}N\parallel_{N}(\mbox{ i.e. }=1),\,\,1\leq i\leq m,\mbox{ and }\ell(\psi(wN))=\parallel wN\parallel_{N}.$
Let $\varphi$ be the homomorphism from $\mathsf{F}$ to $H$ which induces $\psi$ on the free basis of $\mathsf{F}$. It is easy to see that the map $vN\to\varphi(v)$, $v\in F$, induces $\psi$ on $FN$. If $\varphi(w)\in\varphi(B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N)$ then the value $\ell(\varphi(wN))$ is not greater than $m$ by the triangle inequality. This leads to a contradiction with the equality $\parallel wN\parallel_{N}=\ell(\varphi(wN))$.
@@
Take a finite $Q\subset\mathbb{N}$. Let $n=\mathsf{max}(|f|\,|\,f\in F)$. We may assume that $Q$ is an initial segment of the set of natural numbers and contains $\mathsf{max}\{\parallel fN\parallel_{N}\,|\,f\in F\cup B^{\mathsf{F},|\cdot|}_{2n}(1)\}+1$. Note that $n_{f}=\parallel fN\parallel_{N}$ for each $f\in F$. For every $w\in(F\cup B^{\mathsf{F},|\cdot|}_{2n}(1))\setminus\{1\}$ there is a subgroup $H_{w}\triangleleft\,\mathsf{F}$ of finite index such that
  • the map $fN\to fH_{w}$ is a partial isomorphism on $\{fN|f\in(F\cup B^{\mathsf{F},|\cdot|}_{2n}(1))\}$,
@@
Now it is easy to find a subgroup of finite index of $\mathsf{F}$, say $H_{F}$, such that these conditions are satisfied simultaneously for all $w\in(F\cup B^{\mathsf{F},|\cdot|}_{2n}(1))\setminus\{1\}$ and $m\in Q$ with $H_{F}$ instead of $H_{w}$.
@@
Assume that the set $S\subseteq\mathsf{F}$ defining the word metric $\parallel\cdot\parallel_{N}$ is finite and that $N<\mathsf{F}$ is finitely generated as a normal subgroup. Then the condition that the metric group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically LEF implies that it is metrically RF.
@@
Assume that $N$ is normally generated by words $w_{1},\ldots,w_{k}\in\mathsf{F}$. Let $m>\mathsf{max}(|w_{1}|,\ldots,|w_{k}|)$. If a finite set $F\subset\mathsf{F}$ contains all words of length $\leq m$ then $N$ is included into the kernel of every homomorphism $\varphi$ from $\mathsf{F}$ onto a finite group $H$ such that the corresponding map $fN\to\varphi(f)$, $f\in F$, is a partial isomorphism $\mathsf{F}/N\to H$ defined on $\{fN|f\in F\}$. In particular such a $\varphi$ induces a homomorphism $\mathsf{F}/N\to H$. We now see that the condition of Theorem 5.1 implies that each $B^{\mathsf{F},\parallel\cdot\parallel}_{n}(1)N$ is closed in the profinite topology of $\mathsf{F}$. By Theorem 3.3 we obtain the conclusion of this corollary. ∎
@@ 5.2 Connections with metric weak soficity
As we already know metrically LEF groups are metrically weakly sofic. When does the converse hold? In order to answer this question we apply a kind of flexible stability, [2]. Note that a metric group $(G,d)$ with $d\leq 1$ is metrically weakly sofic if and only if it isometrically embeds into a metric ultraproduct of finite metric groups. In the following definition we give additional restrictions on such embedding.
@@
We say that a metric group $(G,\ell)$ is LEF (flexibly) stable if for every finite subset $F\subset G$ there is a function $f_{F}:[0,+\infty)\rightarrow[0,+\infty)$ with $\lim_{x\to 0}f(x)=0$ such that for all $\delta$, $(C,\ell_{C})$ and $\phi$ satisfying the conditions of Definition 1.3 there is a finite normed group $(D,\ell_{D})$ and an injective map $\psi:F\rightarrow D$ such that any triple $h,g,hg\in F$ satisfies $\psi(hg)=\psi(h)\psi(g)$ and any $g\in F$ satisfies $|\ell(g)-\ell_{D}(\psi(g))|<f_{F}(\delta)$.
@@
Proof. Let us fix a free base of $\mathsf{F_{n}}$ which can be infinite. Let $F\subset\mathsf{F_{n}}$ be finite. Find $\bar{b}$, the minimal initial segment of the basis such that $F$ belongs to some ball $B_{k}(1)$ with respect to $\bar{b}$. Take the minimal $k$ with this property, i.e. $\mathsf{diam}(F)=2k$ with respect to $\bar{b}$. We will prove that the function $f_{F}(x)=3k\cdot x$ satisfies Definition 5.3. We may assume that $\bar{b}\subseteq F$ and $F$ is closed under taking subwords.
Let $\delta>0$ and $\phi:F\rightarrow C$ be a map into a finite metric group $(C,\ell_{C})$ as in Definition 1.3. We claim that the group $(C,\ell_{C})$ can be taken as $(D,\ell_{D})$ in Definition 5.3 where $f_{F}$ is as above. Indeed, since $\mathsf{F_{n}}$ is free the map $\bar{b}\rightarrow\phi(\bar{b})$ extends to a homomorphism $\mathsf{F_{n}}\rightarrow C$. We call it $\psi$. When $b_{i}$ and $b_{j}$ are taken from $\bar{b}$ then
@@
and $d_{C}(\psi(b_{i}b_{j}),\phi(b_{i}b_{j}))\leq\delta$ (assuming that $b_{i}b_{j}\in F$). One can similarly show that $d_{C}(\phi(b_{i}^{-1}),\psi(b_{i}^{-1}))\leq\delta$. Using invariantness of $d_{C}$ and properties of $\phi$ we obtain
$d_{C}(\phi(b^{\varepsilon_{1}}_{i}b^{\varepsilon_{2}}_{j}),\psi(b^{\varepsilon_{1}}_{i}b^{\varepsilon_{2}}_{j}))=d_{C}(\phi(b^{\varepsilon_{1}}_{i}b^{\varepsilon_{2}}_{j}),\psi(b_{i})^{\varepsilon_{1}}\psi(b_{j})^{\varepsilon_{2}})\leq 3\delta,$
Applying this argument inductively we obtain that when $v\in F$ and $|v|=\ell$ then
$d_{C}(\phi(v),\psi(v))<2\ell\cdot\delta,$
@@
$|\ell_{C}(\phi(v))-\ell_{C}(\psi(v))|<2\ell\cdot\delta.$
On the other hand $|\parallel v\parallel-\ell_{C}(\phi(v)|<\delta$. In particular
$|\parallel v\parallel-\ell_{C}(\psi(v)|<3\ell\cdot\delta\leq 3k\cdot\delta.$
@@
It is proved in Lemma 5 of [15] that connected abelian Lie groups with the “eulidean” length function are metrically weakly sofic. The proof given in that paper shows that

### v2

@@ 3.3 When metric RF does not depend on metric
Assume that $G$ is normally generated by finite $S_{1}$ and $S_{2}$. Let $\parallel\cdot\parallel_{1}$ and $\parallel\cdot\parallel_{2}$ be the corresponding word metrics of $G$. The following question looks very interesting. Does the property that $G$ is metrically LEF (metrically RF or metrically weakly sofic) with respect to one of the norms $\parallel\cdot\parallel_{1}$ and $\parallel\cdot\parallel_{2}$ imply the same property for another norm? We do not even know if these word norms generate the same topology in the following sense. Note that the topology defined by $\parallel\cdot\parallel_{i}$, $i\in\{1,2\}$, according to the recipe of Section 2.1, is the discrete $\{0,1\}$-topology. On the other hand if we exclude from the sub-base of all $B_{<r}(g)$ the balls of size 1 then we obtain a coarser topology which is not necessarily discrete. We will say that this is the corrected topology defined by the norm. Using Theorem 3.4 we see the following statement.
@@ 4 Metric weak soficity
Assume that the set $S\subseteq\mathsf{F}$ defining the word norm $\parallel\cdot\parallel_{N}$ is finite. The normed group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically weakly sofic if and only if for any finite sequence $g_{1}$, $g_{2}$,…, $g_{k}\in B^{\mathsf{F},\parallel\cdot\parallel}_{1}(1)$ and $h_{1}$, $h_{2}$,…, $h_{l}$ from $N$ (possibly with repetitions) the closure of every product $g_{i_{1}}^{\mathsf{F}}\cdot g_{i_{2}}^{\mathsf{F}}\ldots\cdot g_{i_{m}}^{\mathsf{F}}\cdot h_{1}^{\mathsf{F}}\cdot h_{2}^{\mathsf{F}}\ldots h_{l}^{\mathsf{F}}$ is a subset of $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$.
@@
Necessity. We want to show that the complement of $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ does not intersect the closure of the product $g_{i_{1}}^{\mathsf{F}}\cdot g_{i_{2}}^{\mathsf{F}}\ldots\cdot g_{i_{m}}^{\mathsf{F}}\cdot h_{1}^{\mathsf{F}}\cdot h_{2}^{\mathsf{F}}\ldots h_{l}^{\mathsf{F}}$. Assume $w\not\in B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$. Let $s=\parallel wN\parallel_{N}$. Thus $s>m$. Let $n$ be a natural number such that
@@
Let $\tilde{\phi}$ be the homomorphism $\mathsf{F}\to H$ defined by $\phi$ on the free basis of $\mathsf{F}$. It is easy to see that for any $v\in B^{\mathsf{F},|\cdot]}_{<n}(1)$ we have $\ell((\tilde{\phi}(v))^{-1}\phi(vN))<2n\varepsilon$ (we should take double $n\varepsilon$ since inverses of the elements of the base can also appear in $v$). It suffices to show that $\tilde{\phi}(w)$ does not belong to
@@
We view $\phi$ on cosets from $B^{\mathsf{F},|\cdot|}_{n}(1)N$ as a map which assigns $\phi(f)$ to $fN$ with $f\in B^{\mathsf{F},|\cdot|}_{n}(1)$. In order to verify the weak soficity condition for $\phi$, $\varepsilon$ and elements of $FN$ note that when $f_{1}N\cdot f_{2}N=f_{3}N$ for $f_{1},f_{2},f_{3}\in F$ we have that $\phi(f_{1})\phi(f_{2})(\phi(f_{3}))^{-1}$ is equal to some $\phi(h)$ with $h\in N$, i.e an element of the norm $\varepsilon$.
@@
In the assumptions of the theorem admit that $S$ can be infinite and can meet infinitely many conjugacy classes in $\mathsf{F}$. Then if the normed group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically weakly sofic then for any finite sequence $g_{1}$, $g_{2}$,…, $g_{k}\in B^{\mathsf{F},\parallel\cdot\parallel}_{1}(1)$ and $h_{1}$, $h_{2}$,…, $h_{l}$ from $N$ (possibly with repetitions) the closure of the product $g_{i_{1}}^{\mathsf{F}}\cdot g_{i_{2}}^{\mathsf{F}}\ldots\cdot g_{i_{m}}^{\mathsf{F}}\cdot h_{1}^{\mathsf{F}}\cdot h_{2}^{\mathsf{F}}\ldots h_{l}^{\mathsf{F}}$ is a subset of $B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$.
@@
The normed group $(\mathsf{F}_{2},\parallel\cdot\parallel)$ is not metrically weakly sofic.
@@ 5.1 Characterization of metric LEF
Our characterization of metric LEF clarifies the situations of Sections 3 and 4. We preserve the notation of Section 3.1.
@@
Assume that the set $S\subseteq\mathsf{F}$ defining the word metric $\parallel\cdot\parallel_{N}$ is finite. The normed group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically LEF if and only if for any finite set $F\subset\mathsf{F}$ representing pairwise distinct elements modulo $N$, for any natural number $m$ and any $w\in\mathsf{F}\setminus B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ there is a homomorphism $\varphi$ from $\mathsf{F}$ onto a finite group $H$ such that $\varphi(w)\not\in\varphi(B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N)$ and the corresponding map $fN\to\varphi(f)$, $f\in F$, is a partial isomorphism $\mathsf{F}/N\to H$ defined on $\{fN|f\in F\}$.
@@
Necessity. Take any finite $F$, a number $m$ and $w\in\mathsf{F}\setminus B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N$ as in the formulation of the theorem. We may assume that $F$ contains $w$, a list of representatives of the conjugacy classes of $S^{\mathsf{F}}$, and the free generators of $\mathsf{F}$, and, furthermore, it is closed under the operation of taking subwords of a word.
Let $n$ be a natural number such that
$|F|(\mathsf{max}(|f|\,|\,f\in F))+|S|(\mathsf{max}(|g|\,|g\in S))+|w|<n,$
Since $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically LEF, there is a finite normed group $(H,\ell)$ and a map $\psi:{\mathsf{F}}/N\rightarrow H$ witnessing metric LEF for all words from $B^{\mathsf{F},|\cdot|}_{<n}(1)$. According to Remark 2.8 and Proposition 2.10 we may assume that $\ell$ is integer-valued and the norm inequalities in the definition of LEF are equalities. In particular,
[…]
Let $\varphi$ be the homomorphism $\mathsf{F}\to H$ defined by $\psi$ on the free basis of $\mathsf{F}$. It is easy to see that the map $vN\to\varphi(v)$, $v\in B^{\mathsf{F},|\cdot|}_{<n}(1)$, induces $\psi$ on $B^{\mathsf{F},|\cdot|}_{<n}(1)N$.
If $\varphi(w)\in\varphi(B^{\mathsf{F},\parallel\cdot\parallel}_{m}(1)N)$ then the value $\ell(\varphi(w))$ is not greater than $m$ by the triangle inequality. This leads to a contradiction with the equality $\parallel wN\parallel_{N}=\ell(\varphi(w))$.
@@
Take a finite $Q\subset\mathbb{N}$. Let $n=\mathsf{max}(|f|\,|\,f\in F)$. We may assume that $Q$ is an initial segment of the set of natural numbers and contains $\mathsf{max}\{\parallel fN\parallel_{N}\,|\,f\in B^{\mathsf{F},|\cdot|}_{2n}(1)\}+1$. Note that $n_{f}=\parallel fN\parallel_{N}$ for each $f\in F$. For every $w\in B^{\mathsf{F},|\cdot|}_{2n}(1)\setminus\{1\}$ there is a subgroup $H_{w}\triangleleft\,\mathsf{F}$ of finite index such that
  • the map $fN\to fH_{w}$ is a partial isomorphism on $\{fN|f\in B^{\mathsf{F},|\cdot|}_{2n}(1)\}$,
@@
Now it is easy to find a subgroup of finite index of $\mathsf{F}$, say $H_{F}$, such that these conditions are satisfied simultaneously for all $w\in B^{\mathsf{F},|\cdot|}_{2n}(1)\setminus\{1\}$ and $m\in Q$ with $H_{F}$ instead of $H_{w}$.
@@
Assume that the set $S\subseteq\mathsf{F}$ defining the word metric $\parallel\cdot\parallel_{N}$ is finite and that $N<\mathsf{F}$ is finitely generated as a normal subgroup. Then the condition that the normed group $(\mathsf{F}/N,\parallel\cdot\parallel_{N})$ is metrically LEF implies that it is metrically RF.
@@
Assume that $N$ is normally generated by words $w_{1},\ldots,w_{k}\in\mathsf{F}$. Let $m>\mathsf{max}(|w_{1}|,\ldots,|w_{k}|)$. If a finite set $F\subset\mathsf{F}$ contains all words of length $\leq m$ then $N$ is included into the kernel of every homomorphism $\varphi$ from $\mathsf{F}$ onto a finite group $H$ such that the corresponding map $fN\to\varphi(f)$, $f\in F$, is a partial isomorphism $\mathsf{F}/N\to H$ defined on $\{fN|f\in F\}$. In particular such a $\varphi$ induces a homomorphism $G=\mathsf{F}/N\to H$. We now see that the condition of Theorem 5.1 implies that each $B^{G}_{n}(1)$ is closed in the profinite topology of $G$. By Theorem 3.4 we obtain the conclusion of this corollary. ∎
@@ 5.2 Connections with metric weak soficity
As we already know metrically LEF groups are metrically weakly sofic. When does the converse hold? In order to answer this question we apply a kind of flexible stability, [2]. Note that a metric group $(G,d)$ with $d\leq 1$ is metrically weakly sofic if and only if it isometrically embeds into a metric ultraproduct of finite metric groups. In the following definition we give additional restrictions on such an embedding.
@@
We say that a normed group $(G,\ell)$ is LEF (flexibly) stable if for every finite subset $F\subset G$ there is a function $f_{F}:[0,+\infty)\rightarrow[0,+\infty)$ with $\lim_{x\to 0}f(x)=0$ such that for all $\varepsilon$, $(C,\ell_{C})$ and $\phi$ satisfying the conditions of Definition 1.3 there is a finite normed group $(D,\ell_{D})$ and an injective map $\psi:F\rightarrow D$ such that any triple $h,g,hg\in F$ satisfies $\psi(hg)=\psi(h)\psi(g)$ and any $g\in F$ satisfies $|\ell(g)-\ell_{D}(\psi(g))|<f_{F}(\varepsilon)$.
@@
Proof. Let us fix a free base of $\mathsf{F_{n}}$ (which can be infinite). Let $F\subset\mathsf{F_{n}}$ be finite. Find $\bar{b}$, the minimal initial segment of the basis such that $F$ belongs to some ball $B_{k}(1)$ with respect to $\bar{b}$. Take the minimal $k$ with this property, i.e. $\mathsf{diam}(F)=2k$ with respect to $\bar{b}$. We will prove that the function $f_{F}(x)=3k\cdot x$ satisfies Definition 5.3. We may assume that $\bar{b}\subseteq F$ and $F$ is closed under taking subwords.
Let $\varepsilon>0$ and $\phi:F\rightarrow C$ be a map into a finite metric group $(C,\ell_{C})$ as in Definition 1.3. We claim that the group $(C,\ell_{C})$ can be taken as $(D,\ell_{D})$ in Definition 5.3 where $f_{F}$ is as above. Indeed, since $\mathsf{F_{n}}$ is free, the map $\bar{b}\rightarrow\phi(\bar{b})$ extends to a homomorphism $\mathsf{F_{n}}\rightarrow C$. We call it $\psi$. When $b_{i}$ and $b_{j}$ are taken from $\bar{b}$ then
@@
and $d_{C}(\psi(b_{i}b_{j}),\phi(b_{i}b_{j}))\leq\varepsilon$ (assuming that $b_{i}b_{j}\in F$). One can similarly show that $d_{C}(\phi(b_{i}^{-1}),\psi(b_{i}^{-1}))\leq\varepsilon$. Using invariantness of $d_{C}$ and properties of $\phi$ we obtain
$d_{C}(\phi(b^{\epsilon_{1}}_{i}b^{\epsilon_{2}}_{j}),\psi(b^{\epsilon_{1}}_{i}b^{\epsilon_{2}}_{j}))=d_{C}(\phi(b^{\epsilon_{1}}_{i}b^{\epsilon_{2}}_{j}),\psi(b_{i})^{\epsilon_{1}}\psi(b_{j})^{\epsilon_{2}})\leq 3\varepsilon,$
Applying this argument inductively we obtain that when $v\in F$ and $|v|=t$ then
$d_{C}(\phi(v),\psi(v))<2t\cdot\varepsilon,$
@@
$|\ell_{C}(\phi(v))-\ell_{C}(\psi(v))|<2t\cdot\varepsilon.$
On the other hand $|\parallel v\parallel-\,\ell_{C}(\phi(v)|<\varepsilon$. In particular,
$|\parallel v\parallel-\,\ell_{C}(\psi(v)|<3t\cdot\varepsilon\leq 3k\cdot\varepsilon.$
@@
It is proved in Lemma 5 of [14] that connected abelian Lie groups with the “euclidean” length function are metrically weakly sofic. The proof given in that paper shows that